OpenStudy (ayyookyndall):

What is the sum of the interior angle measures of a 37-gon? Explain how you calculate the answer.

OpenStudy (solomonzelman):

The sum of interior angles with number of sides \(\large\color{slate}{ n }\) (of course \(\large\color{slate}{ n }\) is integer, and \(\large\color{slate}{n\ge3 }\) ) \(\large\color{slate}{ sum~=~180\times(n-2) }\)

OpenStudy (studygurl14):

Sum of the interior angle measures of an n-gon: 180(n-2)

OpenStudy (solomonzelman):

in your case n is 37, as you have probably figured out already.

OpenStudy (ayyookyndall):

Okay now I am confused... @StudyGurl14 @SolomonZelman because where did did you get the two? and whats n?

OpenStudy (solomonzelman):

just fill in the numbers into the formula: if you want a proof for this, then https://answers.yahoo.com/question/index?qid=20080505202850AAe8NIS or use other google souces

OpenStudy (ayyookyndall):

so sum = 180 x (n - 2) sum = 180 x (37 - 2) sum = 180 x (35) sum = 6300 am i correct? @SolomonZelman

OpenStudy (solomonzelman):

yes, that is totally correct.